OpenStudy (anonymous):

Consider the line that passes through the points (2, 3) and (6, 11). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y -11 = 2(x - 6) y - 3 = 2(x - 2) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the t

OpenStudy (anonymous):

Neither equation #1 nor equation #2 represents the line that passes through the two given points.

OpenStudy (saifoo.khan):

How did you got that?

OpenStudy (anonymous):

school

OpenStudy (anonymous):

The given points (x1,y1) and (x2,y2) are (2,3) and (6,11) Substituting these two points in equation 1 and 2: We get that the equation 1 does not pass through (2,3) since the equation is not reduced to 0 on substitution. Where as the equation 1 passes through the point (6,11) since the equation reduces to 0 on substitution. Similarly we substitute point (6,11) and find that only equation 2 passes through it. Hence we find that "Neither equation #1 nor equation #2 represents the line that passes through the two given points."

OpenStudy (anonymous):

thats what i got

OpenStudy (saifoo.khan):

@Snapbacklive , lol. im not an Arab.

OpenStudy (anonymous):

i know, just thought it would be funny.

OpenStudy (saifoo.khan):

points (2, 3) and (6, 11). y -11 = 2(x - 6) 3 - 11 = 2(3-6) -8 = 2(-3) -8 = -6 So, it's NOT a solution for line 1.

OpenStudy (saifoo.khan):

Line 2: y - 3 = 2(x - 2) points (2, 3) and (6, 11). 3 -3 = 2(2-2) 0 = 0 Using second point: (6,11) 11 - 3 = 2(6-2) 8 = 2(4) 8 = 8 So it is a solution! PROVED.

OpenStudy (saifoo.khan):

@Snapbacklive , hehe yup! It's funny.

OpenStudy (anonymous):

yeah